If the consumption function is given as C= 20 + 0.8Yd, complete the 

 Following table: 

 

Disposable 

Income (Yd) 

200 500 

Consumption 

(C) 

 300 700 

Savings (S) 100


1
Expert's answer
2021-03-09T07:25:24-0500
"solution"

The "C=20+0.8Yd"


"given\ disposable \ income(Yd) :\\
at \ 200\\
C=20+(0.8\times 200)=180\\
at\ 500\\
C=20+(0.8\times500)=420\\"


"Given\ Consumption (C):\\
at \ 300\\
300=20+0.8Yd=350\\
700=20+0.8Yd=850\\"


"Savings(S)\\
S=Y-C\\
100=Y-20+0.8Yd\\
120=Y+0.8Yd\\
120=1.8Yd\\
Y=[\frac{120}{1.8}]\\
Y=66.66"